OpenStudy (anonymous):

f(x)=ln(x)+e^(1/x-1). What´s the domain for f? Any suggestions on where to start? :)

OpenStudy (anonymous):

since the function f(x) has two distinct parts, you should start with them and then find the intersection of the two individual domains, like for Ln(x) the domain would be_

OpenStudy (amistre64):

*union of them

OpenStudy (zzr0ck3r):

You have two things to worry about \(x>0\) because of the domain of \(\ln(x)\) and \(x-1\ne 0\) because of the \(\dfrac{1}{x-1}\) (we cant divide by \(0\)). So we can use all positive numbers except \(1\). Domain is \(\{x\mid x>0\text{ and } x\ne 1\}=(0,1)\cup (1,\infty)\) Make sense?

OpenStudy (amistre64):

hmm, might be looking at it in a mirror tho ... the union of the bad parts that is

OpenStudy (anonymous):

Yes, it makes totally sense! Thanks!

OpenStudy (zzr0ck3r):

np

OpenStudy (zzr0ck3r):

I think we want the intersection of the two @amistre64

OpenStudy (anonymous):

np? :)

OpenStudy (zzr0ck3r):

no problem @Wikis

OpenStudy (anonymous):

Of course XD Thought it was a math term that I had missed :D

OpenStudy (amistre64):

intersection of the domains yes ... i was thinking about the exclusions at first. we want to make sure all the bad parts are avoided for each term

OpenStudy (amistre64):

if x=2,3 are bad for the first part, and x=2,4,5 are bad for the last part; then we would exclude x=2,3,4,5 was in my head

OpenStudy (amistre64):

time to build a house ... have fun